OpenStudy (anonymous):

What kind of system is the set of equations below? 2x - 3y = 3 -4x + y = -6

OpenStudy (anonymous):

Can Someone please tell me the answer for this one

OpenStudy (anonymous):

2x-3y=3 -4x+y=-6 simt ---method 16x+12y=-12 16x+4y=-12 16y=-36 y=2 then2x-3y=2 but y=2 2x-6=2 2x=8 x=4 y=2and x=4

OpenStudy (anonymous):

What Kind Of System is it?

OpenStudy (anonymous):

is it consistent=a single solution??